y=7-5(x+4)
Solve in intercept form(show your work algebraically)

Answer by jim_thompson5910(35256)

y=7-5(x+4)

y=7-5(x)-5(4)

y = 7-5x-20

y = -5x - 13


So the answer is y = -5x - 13